Evaluate the integral of (sin³ x + sinxcos²x)dx


1
Expert's answer
2021-04-14T13:01:10-0400

We have to find "\int" (sin³(x) + sin(x)cos²(x)) dx

Let I = "\int" (sin³(x) + sin(x)cos²(x)) dx

= "\int" sin(x)(sin2(x) + cos²(x)) dx

= "\int" sin(x) dx "\because" sin2(x) + cos²(x) = 1

= - cos(x) + C where C is a constant

hence,

I = - cos(x) + C